If {c,d} is the solution set of x^2-7=6x evaluate the following: c^3+c^2d+cd^2+d^3
Answer by edjones(8007)

x^2-6x-7=0
(x-7)(x+1)=0
x=7 x=-1
so (c,d)=(7,-1)
c^3+c^2d+cd^2+d^3
=7^3+7^2*-1+7(-1)^2+(-1)^3
=343-49+49-1=342
Ed